# DCT-CryptoNets: Scaling Private Inference in the Frequency Domain

### v1

@@ 4.3 Evaluation of Accuracy and Latency
CryptoDL [12] CIFAR-10 3$\times$322 Custom CNN BGV 91.4% 11,686 $\sim$
Faster Cryptonets [7] CIFAR-10 3$\times$322 Custom CNN FV-RNS 75.9% 3,240 $\sim$
SHE [21] CIFAR-10 3$\times$322 Custom CNN TFHE 92.5% 2,258 470
Lee et al. (1) [19] CIFAR-10 3$\times$322 ResNet-20 CKKS 92.4% 10,602 $\sim$
Lee et al. (2) [18] CIFAR-10 3$\times$322 ResNet-20 CKKS 91.3% 2,271 $\sim$
Kim et al. [16] CIFAR-10 3$\times$322 Plain-20 CKKS 92.1% 368 $\sim$
Ran et al. [29] CIFAR-10 3$\times$322 ResNet-20 CKKS 90.2% 392 $\sim$
Rovida et al. [31] CIFAR-10 3$\times$322 ResNet-20 CKKS 91.7% 336 $\sim$
Benamira et al. [1] CIFAR-10 3$\times$322 VGG-9 TFHE 74.0% 570 48
Stoian et al. [32] CIFAR-10 3$\times$322 VGG-9 TFHE 87.5% 18,000 3,000
@@
SHE [21] CIFAR-10 3$\times$322 ResNet-18 TFHE 94.6% 12,041 2,509
@@
SHE [21] ImageNet 3$\times$2242 ResNet-18 TFHE 66.8% 216,000 45,000
@@
DCT-CryptoNets demonstrates competitive accuracy when compared to existing methods on both CIFAR-10 and ImageNet (see Table 4). While CKKS-based methods demonstrate slightly lower latency on smaller models due to their utilization of single instruction multiple data (SIMD) packing and polynomial approximations of activation functions (PAAs), the inherent limitations of PAAs hinder CKKS-based methods from scaling to larger networks and high-resolution images like those in ImageNet. It’s crucial to note that CKKS’s focus on amortized runtime through SIMD packing, a feature unavailable in TFHE, makes direct latency comparisons via normalization less meaningful.
Comparing DCT-CryptoNets to SHE [21], the only other method demonstrating scalability to ImageNet, we observe comparable accuracy and latency for smaller models and images. It’s worth noting that SHE’s fastest model on CIFAR-10 is a custom CNN without skip connections, contributing to its slight latency advantage over our ResNet-20 model. As network or image size increases, DCT-CryptoNets showcases superior scalability, with progressively greater latency improvements compared to SHE culminating in a 5.3$\times$ speedup on ImageNet. Furthermore, even without DCT optimizations, larger DCT-CryptoNets ResNet-18 models outperform SHE on both CIFAR-10 and ImageNet. This improvement can be attributed to SHE’s reliance on a leveled TFHE scheme, which necessitates higher multiplicative depths and larger ciphertexts leading to latency amplifications in deeper networks. In contrast, by leveraging programmable bootstrapping, DCT-CryptoNets mitigates these limitations and enables efficient handling of larger models.
Overall, our results demonstrate that DCT-CryptoNets achieves performance on par with previous work on smaller images and networks, while showcasing superior scalability and efficiency gains when applied to larger, more complex tasks.
@@ 4.4 Limitations and Future Directions
A key limitation of our approach is that performance is highly dependent on the cryptographic hyperparameters and the degree of bit precision during quantization-aware training (QAT). Careful tuning of precision rounding, programmable boostrapping (PBS) error probability and QAT bit-level is essential to navigate the trade-off between accuracy and latency. Future steps should be made to improve the cryptographic and quantization parameter selection process.
Furthermore, the accuracy benefits from DCT optimizations are less pronounced on smaller images (e.g., 32$\times$32) compared to larger ones. This observation aligns with the inherent nature of small images, which often contain less high-frequency information that DCT filtering typically removes. Consequently, the impact of focusing on low-frequency components is less significant for these smaller inputs. Future work could explore techniques to enhance the accuracy benefits of DCT for smaller images, potentially through adaptive filtering or alternative frequency-domain representations.
@@ 5 Conclusion
In this paper, we presented DCT-CryptoNets, a novel framework that addresses the critical challenges of scalability and computational cost in fully homomorphic encrypted neural networks (FHENNs), particularly for high resolution images. By combining frequency-domain learning through the Discrete Cosine Transform (DCT), DCT-CryptoNets significantly reduces the computational burden of non-linear activations and homomorphic bootstraps, enabling more efficient inference on large-scale datasets like ImageNet. Our approach demonstrates substantial latency improvements, achieving a 5.3$\times$ speedup compared to the only prior work on ImageNet and a boost in the reliability of encrypted accuracy variability. Furthermore, isolating the impact of DCT-based optimizations reveals a latency improvement ranging from 1.7$\times$ to 2.4$\times$. These advancements showcase the promise of frequency-domain techniques for privacy-preserving deep learning. This not only represents a significant step towards practical deployment of FHE in real-world applications, but also opens new avenues for future research in optimizing private inference for deep learning.
@@ Acknowledgements
This project was supported from the Purdue Center for Secure Microelectronics Ecosystem – CSME#210205.
@@ Appendix A Appendix
A.1 Quantization Ablation on ImageNet
@@
To examine the impact of quantization on performance, an ablation study was performed varying quantization levels on ImageNet, as detailed in Table 5. Adopting 5-bit quantization, consistent with the approach in SHE [21], resulted in a notable accuracy improvement compared to 4-bit quantization, with a manageable increase in latency of approximately 25%. Further increasing the quantization level to 8-bit yielded only marginal accuracy gains, while incurring a substantial increase in latency. These results suggest that 5-bit quantization strikes an effective balance between accuracy and efficiency for large-scale datasets like ImageNet.
A.2 ReLU Pruning Techniques
$\Delta$: RGB $\rightarrow$ DCT DCT-Net [35] + First Layer Pruned [11] + Skip Connections Pruned [11] + Both Pruned
Ops. #ReLUs -26.1% -34.8% -58.7% -67.4%
#HOPs -21.7% -33.1% -24.3% -25.7%
Acc. CIFAR-10 +0.0% -0.4% -2.67% -2.54%
ImageNet +2.2% +2.3% $\sim$ $\sim$
DCT-Net [35] inherently reduces ReLU operations due to spatial dimension reduction. DCT-CryptoNets further optimizes latency by pruning the first layer’s ReLU’s, with minimal impact on accuracy. However, skip connection ReLU pruning is not recommended as it incurs significant accuracy degradation and increases homomorphic operations even though the number of ReLU operations decreases. This is because removing ReLU after the addition of two quantized tensors necessitates the insertion of a quantization identity function.
A.3 DCT-CryptoNets for Small(er) Images and Networks
When applying DCT optimizations to smaller images and networks (e.g., ResNet-20 on 32$\times$32 images), adaptations are necessary to address the excessive spatial dimension reduction caused by the standard 8$\times$8 DCT filter. We mitigate this by reducing the DCT filter size to 4$\times$4, resulting in 48 frequency channels (16 per YCrCb component) instead of the typical 192. To accommodate this change, we introduce a modified ResNet-20 architecture with two key adjustments (see Fig. 3):
Although this channel expansion doubles the number of parameters, it enables effective DCT utilization on smaller images. As demonstrated in Table 2, DCT-CryptoNets ResNet-20 achieves a 4% accuracy improvement over the traditional ResNet-20 while maintaining comparable latency.
Figure 3: Traditional vs. DCT-CryptoNets ResNet-20 Architecture
A.4 Sensitivity of Cryptographic Hyperparameters in FHENN
Table 7: Comparing the change of latency, top-1 accuracy and ciphertext memory requirements of a 3$\times$2242 model on 1000 samples from ImageNet when modulating two encryption hyperparameters (precision rounding and PBS error probability).
@@ A.4 Sensitivity of Cryptographic Hyperparameters in FHENN
A.5 In the context of MLaaS
DCT-CryptoNets leverages the Discrete Cosine Transform (DCT) to operate directly in the frequency domain, aligning with the prevalent use of DCT in image compression standards like JPEG. This enables seamless integration with existing image processing pipelines and facilitates efficient transmission and processing in machine learning as a service (MLaaS) systems and other applications where image data is sent for remote analysis. In this paradigm (Figure 4), a model is trained locally, encrypted, and then deployed in the cloud. During inference, users transform their images into DCT representations (a process often already performed due to JPEG compression) and encrypt them using their private key. The encrypted image is then sent to the cloud-based model for processing, and the resulting encrypted output is returned to the user. To obtain the final prediction, the user decrypts the output of the penultimate layer (prior to the fully-connected layer) and processes the resulting plaintext tensor through a local classifier.
Figure 4: MLaaS System with DCT-CryptoNets.
A.6 TFHE-based FHENN Compute Resources
@@
SHE [21] CIFAR-10 3$\times$322 Custom CNN 92.5% 2,258 20 470
Benamira et al. [1] CIFAR-10 3$\times$322 VGG-9 74.01% 570 8 48
Stoian et al. [32] CIFAR-10 3$\times$322 VGG-9 87.5% 18,000 16 3,000
@@
SHE [21] CIFAR-10 3$\times$322 ResNet-18 94.62% 12,041 20 2,509
@@
SHE [21] ImageNet 3$\times$2242 ResNet-18 69.4% 216,000 20 45,000
@@
Since TFHE prioritizes individual inference latency, we normalize prior work’s latency values to the 96-thread resource capacity of DCT-CryptoNets to ensure a fair comparison. This direct normalization based on CPU-threads is motivated by the fact that the Concrete-ML library utilizes 100% of available CPU resources. We noticed when running DCT-CryptoNets on a 24-thread Intel Core i9-12900KF there was exactly 4$\times$ increase in latency compared to the 96-thread dual AMD Ryzen Threadripper PRO 5965WX. It is important to note that many CKKS-based methods utilize single instruction multiple data (SIMD) packing to optimize amortized runtime, a feature not available in TFHE. Consequently, normalizing latency for CKKS-based methods would not accurately reflect their performance characteristics due to their inherent reliance on batch processing.

### v2

@@ 4.3 Comparative Benchmarking of Accuracy and Latency
Hesamifard et al. (2017) CIFAR-10 3$\times$322 Custom CNN BGV 91.4% 11,686 $\sim$
Chou et al. (2018) CIFAR-10 3$\times$322 Custom CNN FV-RNS 75.9% 3,240 $\sim$
SHE (Lou & Jiang, 2019) CIFAR-10 3$\times$322 Custom CNN TFHE 92.5% 2,258 470
Lee et al. (2022b) CIFAR-10 3$\times$322 ResNet-20 CKKS 92.4% 10,602 $\sim$
Lee et al. (2022a) CIFAR-10 3$\times$322 ResNet-20 CKKS 91.3% 2,271 $\sim$
Kim & Guyot (2023) CIFAR-10 3$\times$322 Plain-20 CKKS 92.1% 368 $\sim$
Ran et al. (2023) CIFAR-10 3$\times$322 ResNet-20 CKKS 90.2% 392 $\sim$
Rovida & Leporati (2024) CIFAR-10 3$\times$322 ResNet-20 CKKS 91.7% 336 $\sim$
Benamira et al. (2023) CIFAR-10 3$\times$322 VGG-9 TFHE 74.0% 570 48
Stoian et al. (2023) CIFAR-10 3$\times$322 VGG-9 TFHE 87.5% 18,000 3,000
@@
SHE (Lou & Jiang, 2019) CIFAR-10 3$\times$322 ResNet-18 TFHE 94.6% 12,041 2,509
@@
SHE (Lou & Jiang, 2019) ImageNet 3$\times$2242 ResNet-18 TFHE 66.8% 216,000 45,000
@@ 4.4 Limitations and Future Directions
Performance is highly dependent on the cryptographic hyperparameters and the degree of bit precision during quantization-aware training (QAT). Careful tuning of precision rounding, programmable boostrapping (PBS) error probability and QAT bit-level is essential to navigate the trade-off between accuracy and latency. Future steps should be made to improve this hyperparameter selection process. Furthermore, the accuracy benefits from DCT are less pronounced on smaller images (e.g., 32$\times$32) compared to larger ones. This observation aligns with the inherent nature of small images, which often contain less high-frequency information that DCT filtering typically removes. Consequently, the impact of focusing on low-frequency components is less significant for these smaller inputs. Future work could explore techniques to enhance the accuracy benefits of DCT for smaller images.
@@ 5 Conclusion
In this paper, we presented DCT-CryptoNets, a novel framework that addresses the critical challenges of computational cost and scalability in fully homomorphic encrypted neural networks (FHENNs). By utilizing frequency-domain learning through the Discrete Cosine Transform (DCT), DCT-CryptoNets significantly reduces the computational burden of non-linear activations and homomorphic bootstraps by focusing on low-frequency information, enabling more efficient inference on large-scale datasets like ImageNet.
DCT-CryptoNets exhibits several key advantages: (1) it achieves significant latency improvements, demonstrating up to a 5.3$\times$ speedup compared to prior work, with DCT-based optimizations contributing a 1.7$\times$ to 2.4$\times$ speedup in isolation; (2) focusing on perceptually relevant low-frequency information through DCT maintains or even boosts accuracy compared to RGB-based networks; (3) it enhances the reliability of predictions by reducing error-accumulating homomorphic bootstrap operations, resulting in more precise predictions (e.g., from ±2.5% to ±1.0% on ImageNet); and (4) it demonstrates superior scalability, achieving greater reductions in homomorphic operations compared to RGB-based implementations as network and image size increases.
These advancements showcase the promise of frequency-domain techniques for privacy-preserving deep learning. This not only represents a significant step towards practical deployment of FHE in real-world applications, but also opens new avenues for future research in optimizing private inference for deep learning.
Reproducibility Statement
We strongly believe in the importance of reproducibility in scientific research. Therefore, we strive for full transparency in our work. To this end regarding explanations in the paper, the Methodology section (Section 3) provides a comprehensive overview of our model topology, training methodology, and encryption process. The Experimental Setup subsection (Section 4.1) offers further details regarding libraries, tools and hardware environments needed to facilitate reproducibility. All training, quantization, and cryptographic hyperparameters are also documented in the Appendix (Section A.7). To further promote reproducibility, our code has been released at https://github.com/ar-roy/dct-cryptonets.
@@ Acknowledgements
This project was supported from the Purdue Center for Secure Microelectronics Ecosystem – CSME#210205. We wish to also acknowledge the insightful discussions with the Zama Concrete-ML team and within the FHE.org community, which enriched our perspective on this research.
@@ Appendix A Appendix
[…]
A.1 Threat Models and Mitigation Strategies
While fully homomorphic encryption (FHE) provides robust security against many traditional attacks, a comprehensive understanding of potential vulnerabilities within the DCT-CryptoNets framework is crucial for ensuring strong privacy guarantees. This analysis must consider the unique security considerations inherent to FHE-based private inference.
  • Server-side security: The server only receives encrypted DCT coefficients, protecting against a wide range of server-side attacks, including those from an honest-but-curious adversary. However, in the case of a malicious server, additional measures like verifiable computation (VC) may be necessary to ensure correct function evaluation (Viand et al., 2023; Atapoor et al., 2024).
  • Communication-channel security: Encrypted communication between the client and server safeguards against eavesdropping. However, secure key transfer mechanisms are crucial to prevent unauthorized access to the decryption key.
  • Client-side security: Data privacy relies heavily on the security of the client device, as this is where the data exists in plaintext. Robust client-side security measures, such as secure key storage and protection against malware, are essential. While FHE itself is secure against various attacks such as chosen/known plaintext attacks and chosen ciphertext attacks (Peikert, 2016), these protections are contingent on the client’s ability to safeguard their own secret key.
A.2 In the context of MLaaS
Figure 3: MLaaS system with DCT-CryptoNets.
DCT-CryptoNets leverages the Discrete Cosine Transform (DCT) to operate directly in the frequency domain, aligning with the prevalent use of DCT in image compression standards like JPEG. This enables seamless integration with existing image processing pipelines and facilitates efficient transmission and processing in machine learning as a service (MLaaS) systems and other applications where image data is sent for remote analysis. In this paradigm (see Figure 3), a model is trained locally, encrypted, and then deployed in the cloud. During inference, users transform their images into DCT representations (a process often already performed due to JPEG compression) and encrypt them using their private key. The encrypted image is then sent to the cloud-based model for processing, and the resulting encrypted output is returned to the user. To obtain the final prediction, the user decrypts the output of the penultimate layer (prior to the fully-connected layer) and processes the resulting plaintext tensor through a local classifier.
In many FHE-based private inference applications, both a secret key and evaluation key are generated. The evaluation key, analogous to a public key specifying the operations of the FHE circuit, allows for the execution of computations on the encrypted model. This approach ensures that only the user possessing the corresponding secret key can decrypt the resulting inference output:
  • Deployment: A trained fully homomorphic encrypted neural network (FHENN) is deployed to the cloud server.
  • User request: User requests the FHENN cryptographic parameters. Once received they generated both secret and evaluation keys locally. Each client would therefore have their own secret key.
  • Key Exchange: User sends the evaluation key along with their encrypted inputs to the server.
  • Inference: Server runs private inference with the users’ evaluation key and encrypted inputs. Then sends encrypted results back to the user.
  • User decryption: User then decrypts the results with their secret key.
A.3 ReLU Pruning Techniques
$\Delta$: RGB $\rightarrow$ DCT DCT-Net (Xu et al., 2020) + First Layer Pruned (Ghodsi et al., 2020) + Skip Connections Pruned (Ghodsi et al., 2020) + Both Pruned
Ops. #ReLUs -26.1% -34.8% -58.7% -67.4%
#HOPs -21.7% -33.1% -24.3% -25.7%
Acc. CIFAR-10 +0.0% -0.4% -2.67% -2.54%
ImageNet +2.2% +2.3% $\sim$ $\sim$
DCT-Net (Xu et al., 2020) inherently reduces ReLU operations due to spatial dimension reduction. DCT-CryptoNets further optimizes latency by pruning the first layer’s ReLU’s, with minimal impact on accuracy. However, skip connection ReLU pruning is not recommended as it incurs significant accuracy degradation and increases homomorphic operations even though the number of ReLU operations decreases. This is because removing ReLU after the addition of two quantized tensors necessitates the insertion of a quantization identity function in TFHE-based networks.
A.4 DCT-CryptoNets for Small(er) Images and Networks
Figure 4: Traditional vs. DCT-CryptoNets ResNet-20 Architecture. Changes between the two architectures are bolded.
When applying DCT optimizations to smaller images and networks (e.g., ResNet-20 on 32$\times$32 images), adaptations are necessary to address the excessive spatial dimension reduction caused by the standard 8$\times$8 DCT filter. We mitigate this by reducing the DCT filter size to 4$\times$4, resulting in 48 frequency channels (16 per YCrCb component) instead of the typical 192. To accommodate this change, we introduce a modified ResNet-20 architecture with two key adjustments (see Figure 4):
Although this channel expansion doubles the number of parameters, it enables effective DCT utilization on smaller images. As demonstrated in Table 2, DCT-CryptoNets ResNet-20 achieves a 4% accuracy improvement over the traditional ResNet-20 while maintaining latency.
A.5 Quantization Ablation on ImageNet
@@
To examine the impact of quantization on performance, we varied the quantization levels on ImageNet, as detailed in Table 6. Adopting 5-bit quantization, consistent with the approach in SHE (Lou & Jiang, 2019), resulted in a notable accuracy improvement compared to 4-bit quantization, with a manageable increase in latency of approximately 25%. Further increasing the quantization level to 8-bit yielded marginal accuracy gains, while incurring a substantial increase in latency. These results suggest that 5-bit quantization strikes an effective balance between accuracy and efficiency for large-scale datasets like ImageNet.
A.6 Sensitivity of Cryptographic Hyperparameters in FHENN
Table 7: Comparing the change of latency, top-1 accuracy and ciphertext memory requirements of a 3$\times$2242 model on 1000 samples from ImageNet when modulating cryptographic hyperparameters (precision rounding and PBS error probability).
@@ A.6 Sensitivity of Cryptographic Hyperparameters in FHENN
A.7 Training, Quantization and Cryptographic Hyperparameters
Table 8: Training, quantization and cryptographic hyperparameters.
[b] Section Hyperparameter CIFAR-10 mini-ImageNet Imagenette ImageNet Training Epochs 60 90 Batch Size 32 256 Optimizer Adam Adam Learning Rate 1e-3 1e-3 Weight Decay 1e-5 1e-5 Gradient Clipping 0.1 0.1 Dropout 0.2 0.2 Scheduler [20, 45] [30, 60] Decay Factor 0.1 0.1 Quantization Weight Bit-Width 4 4 Weight Quantization Protocol Int8WeightPerTensorFloat [^3] Recommended quantization protocols by Concrete-ML for bit-width > 3. Int8WeightPerTensorFloat [^3] Recommended quantization protocols by Concrete-ML for bit-width > 3. Activation Bit-Width 4 4 Activation Quantization Protocol Int8ActPerTensorFloat [^3] Recommended quantization protocols by Concrete-ML for bit-width > 3. Int8ActPerTensorFloat [^3] Recommended quantization protocols by Concrete-ML for bit-width > 3. Cryptographic Number of Bits 5 5 PBS Error Probability 0.01 0.01 Bit-removal Rounding Threshold 6 7
A.8 TFHE-based FHENN Compute Resources
@@
SHE (Lou & Jiang, 2019) CIFAR-10 3$\times$322 Custom CNN 92.5% 2,258 20 470
Benamira et al. (2023) CIFAR-10 3$\times$322 VGG-9 74.01% 570 8 48
Stoian et al. (2023) CIFAR-10 3$\times$322 VGG-9 87.5% 18,000 16 3,000
@@
SHE (Lou & Jiang, 2019) CIFAR-10 3$\times$322 ResNet-18 94.62% 12,041 20 2,509
@@
SHE (Lou & Jiang, 2019) ImageNet 3$\times$2242 ResNet-18 69.4% 216,000 20 45,000
@@
Table 9 presents latency values normalized to the 96-thread computational capacity of our DCT-CryptoNets implementation. This normalization is justified for TFHE-based schemes, which prioritize individual inference latency, as the Concrete-ML library fully utilizes available CPU resources. Our experiments on a 24-thread Intel Core i9-12900KF confirmed a linear relationship between core count and latency, with a 4x increase observed compared to the 96-thread dual AMD Ryzen Threadripper PRO 5965WX. In contrast, CKKS-based methods often leverage single instruction multiple data (SIMD) packing for amortized runtime optimization, a feature not available in TFHE. Consequently, direct latency normalization may not accurately reflect the performance characteristics of CKKS-based approaches due to their inherent reliance on batch processing.